### v1

# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

Abstract
Large language models (LLMs) have demonstrated promise in emulating human-like responses across a wide range of tasks. In this paper, we propose a novel alignment framework that treats LLMs as agent proxies for human survey respondents, affording a cost-effective and steerable solution to two pressing challenges in the social sciences: the rising cost of survey deployment and the growing demographic imbalance in survey response data. Drawing inspiration from the theory of revealed preference, we formulate alignment as a two-stage problem: constructing diverse agent personas called endowments that simulate plausible respondent profiles, and selecting a representative subset to approximate a ground-truth population based on observed data. To implement the paradigm, we introduce P2P, a system that steers LLM agents toward representative behavioral patterns using structured prompt engineering, entropy-based sampling, and regression-based selection. Unlike personalization-heavy approaches, our alignment approach is demographic-agnostic and relies only on aggregate survey results, offering better generalizability and parsimony. Beyond improving data efficiency in social science research, our framework offers a testbed for studying the operationalization of pluralistic alignment. We demonstrate the efficacy of our approach on real-world opinion survey datasets, showing that our aligned agent populations can reproduce aggregate response patterns with high fidelity and exhibit substantial response diversity, even without demographic conditioning.
Keywords Preference modeling  $\cdot$ Value alignment  $\cdot$ Evaluation
1 Introduction
Pollsters are facing mounting challenges in the contemporary landscape, as public willingness to participate in surveys declines across regions and over time. In the United States, telephone response rates have dropped from 36% in the late 1990s to under 6% today Kennedy (2019), while major federal surveys such as the Current Population Survey (CPS) have experienced a 20-point drop since 2015 U.S. Bureau of Labor Statistics (2025). Likewise, across Europe Jabkowski and Cichocki (2025), response rates for national surveys in 36 European countries have fallen 0.8% per year on average between 1999 and 2018. This erosion of participation undermines statistical power and representativeness, particularly because the decline is not uniform across demographic groups. For instance, recent election polling failures have been traced to differential nonresponse: Republicans and independents are increasingly less likely to answer than Democrats Clinton et al. (2022); Parshall (2024). In the social sciences, this is a classic case of sample selection bias, which, if left unchecked, compromises both predictive accuracy and external validity of the survey-based inference. At the same time, the exorbitant cost of conducting large-scale surveys, due to respondent recruitment, compensation, and extensive piloting, places constraints on experimentation and iteration Kennedy (2015). These challenges highlight a pressing need for scalable, flexible, and theoretically grounded alternatives for modeling human preferences.
The epistemic challenge of capturing human preferences, long studied in the social sciences, now finds a structural parallel in large language model (LLM) alignment, where researchers aim to steer model behavior in line with human values and beliefs. Standard approaches, such as supervised fine-tuning (SFT) Tan et al. (2024) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Christiano et al. (2017), train models to reflect normative ideals of helpfulness, harmlessness, or truthfulness. While effective for building general-purpose assistants, these techniques are inherently monolithic: they produce a single dominant behavioral profile, making them ill-suited for modeling pluralistic human populations. Scaling them to reflect multiple value systems requires retraining or fine-tuning for each group, rendering the approach costly, brittle, and inflexible. Consequently, these pipelines struggle with tasks that demand population-level diversity, such as simulating public opinion or generating group-sensitive predictions.
To address this rigidity, alignment researchers have increasingly turned to post-training conditioning techniques that guide model behavior at inference time. A prominent example is the use of structured prompts, variously termed endowments Horton (2023), personas Castricato et al. (2025), or Silicon Samples Argyle et al. (2023), that encode sociodemographic attributes or behavioral dispositions. Prompted this way, LLMs can emulate individual-level responses with considerable fidelity Bisbee et al. (2024), making them attractive for constructing synthetic populations to pilot surveys or test hypotheses. However, this approach remains limited in key respects. First, it depends heavily on human-curated prompts or ground-truth demographic profiles, which may be unavailable or incomplete. Second, it often produces synthetic responses that lack the natural diversity of real populations: LLMs exhibit central-category bias Wang et al. (2024) and generate insufficient inter-agent variation Bisbee et al. (2024). Thus, while endowments reduce the need for retraining, they increase reliance on demographic templates and fall short of capturing true representational diversity.
These limitations reveal a common assumption in alignment: that behavioral fidelity hinges on demographic accuracy. We challenge this view and propose a new theoretical framework grounded in preference reconstruction. Rather than aligning each synthetic agent to a ground-truth demographic profile, we argue that effective emulation requires only the recovery of the aggregate preference structure of the target population. Drawing on revealed preference theory in economics, we treat each agent as a proxy decison-maker whose responses expose latent beliefs or utilities. Crucially, we do not seek to recover the true basis of individual human preferences. Instead, it suffices to identify a functional basis in the latent preference space, a diverse set of proxy agents that span the opinion manifold, and determine appropriate aggregation weights over them. This abstraction is reflected in human societies: demographically distinct populations may exhibit similar collective attitudes, even if their constituent beliefs diverge. Our approach thus enables a more data- and alignment-efficient route to survey emulation, moving beyond identity-conditioned generation toward flexible distributional modeling.
This perspective also clarifies and concretizes the goals of pluralistic alignment Sorensen et al. (2024), which advocates representing a spectrum of human values rather than optimizing for a singular objective. While compelling in principle, the pluralistic alignment literature has remained largely agnostic about how to measure pluralism or evaluate pluralistic systems. Our framework addresses this gap by proposing concrete criteria for measuring variation and fidelity, along with a mechanism to steer agent distributions efficiently under resource constraints so that the aggregate distribution of answers across the agent ensemble matches that of a given target population. In doing so, we operationalize distributional pluralistic alignment through a lens grounded in economic theory and suited for practical survey applications.
Our contributions are as follows:
  • Theoretical foundation. We formalize preference alignment as a two-stage process: constructing a functional basis of proxy agents and determining aggregation weights, drawing on revealed preference theory to recover population-level preferences.
  • Modular alignment system. We implement this theory via a modular system, Prompts to Proxies (P2P), that combines active endowment generation (to promote response diversity) with regression-based aggregation (to estimate preferences without demographic alignment).
  • Pluralistic alignment in practice. We provide the first operational framework for pluralistic alignment, introducing metrics for diversity and fidelity and propose scalable strategies to balance representational coverage and resource efficiency.
  • Bridge to empirical social science. Our framework enables social scientists to simulate and test preference theories by controlling agent endowments and tracking their effects on response distributions. It also supports survey design and can mitigate nonresponse bias, offering a cost-effective complement to traditional data collection.
  • Testbed for alignment research. The modular design provides a controlled environment for evaluating prompting strategies, enabling alignment researchers to explore steerability and value-sensitive modeling at scale.
2 Modeling Preferences via Functional Bases
Preference modeling under conventional lenses
Human preferences are latent, complex, and often inconsistent, posing a persistent challenge for formal modeling. Economists and social scientists have approached this challenge through multiple modeling frameworks. One prominent approach models humans as rational agents who maximize a stable utility function. While elegant in theory and mathematically tractable, these models often struggle to explain empirical anomalies in observed behavior Kahneman and Tversky (1979); Thaler (1999); Kahneman and Tversky (2000). A complementary framework, revealed preference theory Samuelson (1938); Mas-Colell et al. (1995), shifts the focus from explicit utility functions to observable choices. It asks whether a pattern of choices can be rationalized as if generated by a utility-maximizing agent without assuming a specific functional form. This idea has proven especially valuable in empirical settings, offering a way to assess whether behavior is consistent with stable preferences directly from observed data.
We take inspiration from the revealed preference perspective—not in its full rationality axioms, but in its core idea: preferences, though hidden, can be inferred from behaviors, such as responses to survey questions. This principle also underlies modern alignment pipelines such as SFT and RLHF, where models are steered toward human-preferred outputs using behavioral signals in the form of labeled data or feedback on generated responses.
Our framework
To formalize our approach, we zoom out from individual preferences and return to the level of survey analysis. Surveys are not designed to recover each respondent’s utility function; rather, they aim to capture aggregate population preferences, often through summary statistics such as response distributions or marginal means. This shift in perspective from individual to population invites a functional view (Figure 1): we can represent the population preference as a linear combination of individual preference vectors in a latent preference space. Responses to survey questions can then be viewed as mappings from this latent space to an observable response space. Crucially, this mapping is generally many-to-one: different underlying preferences may yield identical observed responses. Yet, as the number and diversity of responses increase, the expressive capacity of the response space improves. In the limit, a sufficiently rich battery of responses enables an increasingly precise identification of latent preferences. This perspective motivates our framework: Instead of recovering the ground-truth preference of each individual, we can construct a functional basis, a diverse set of proxy agents whose preferences span the latent space. By observing their responses and estimating the appropriate aggregation weights, which can be positive or negative, we can reconstruct the aggregate preference structure of the population.
Figure 1: Top: Finding a functional basis to recover latent ground-truth preference. Bottom: Identification of latent preference through multiple observations.
Alignment as a two-stage problem
Based on our theoretical framework, we can cast alignment as a two-stage problem. The first stage involves constructing a broad and expressive candidate basis of proxy agents, each representing a distinct vector in the latent preference space. For the basis to be effective, it must capture meaningful variation across the population. This objective aligns closely with recent calls for pluralistic alignment Sorensen et al. (2024), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal. However, we do not have direct access to individual preferences, and the prompt space is far too vast to explore in an unstructured manner. To address this, we adopt a structured prompting strategy based on attributes—interpretable factors whose variation is likely to influence preference expression. These attributes may be informed from theory (e.g., ideology, values, levels of altruism) or content (e.g., topical relevance). By combining and varying these attributes systematically, we generate a set of endowment-conditioned agents that span different positions in the preference space.
The second stage concentrates on designing an aggregation mechanism to reconstruct the population preference from proxy agents. We frame this as a supervised learning problem: given aggregate responses for a subset of survey questions, we estimate a set of aggregate weights over the proxy agents based on their individual response patterns. These weights are then used to predict responses on held-out questions, enabling out-of-sample evaluation. This setup allows us to examine whether the weighted agentic ensemble indeed captures the target population preference, rather than overfitting to the training set. We adopt regression-based methods for this purpose, as they align naturally with our assumption that population preferences emerge as linear combinations of individual vectors in the latent space. We describe implementation details, including attribute construction, endowment generation, and constrained regression methods, in the following section.
3 Prompts to Proxies (P2P)
To implement our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2). The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained variable section algorithms form the backbone. They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate the results of observed population-level surveys.
Figure 2: System Overview of the Prompts to Proxies (P2P) Framework. P2P operationalizes preference alignment as a two-stage process inspired by revealed preference theory. In Stage 1, active endowment generation uses a structured attribute bank and entropy-guided adaptive sampling strategies to construct diverse agent personas, called endowments. In Stage 2, regression-based aggregation assigns weights to reconstruct population-level preference patterns. An external validity check accesses how well the weighted ensemble predicts aggregate responses on held-out questions, assessing out-of-sample generalization.
3.1 Survey Splitting and Evaluation Protocol
To support the alignment procedure, we partition the full set of survey questions, denoted $\mathcal{Q}$, into three disjoint subsets: $\mathcal{Q}_{\mathrm{train}}$, $\mathcal{Q}_{\mathrm{valid}}$, and $\mathcal{Q}_{\mathrm{test}}$, corresponding to the training, validation, and test sets, respectively. Let $N=|\mathcal{Q}|$ denote the total number of questions, and $N_{\mathrm{train}}$, $N_{\mathrm{valid}}$, and $N_{\mathrm{test}}$ the cardinalities of the three subsets. P2P allows users to set the split sets and ratio flexibly.
For the alignment system, $\mathcal{Q}_{\mathrm{train}}$ is used for active endowment generation. $\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in regression-based aggregation. We reserve $\mathcal{Q}_{\mathrm{test}}$ as an ultimate test of the predictive accuracy of the selected agent ensemble, ensuring that the reconstructed preference has a decent degree of generalizability. For regression analysis, categorical questions are binarized into question-option pairs.
3.2 Attributes as Control Handles
As a crucial part of the active endowment generation logic, we use attributes to shape model preferences. This approach, though counter-intuitive at first glance, is in fact a deliberate reverse engineering design choice. In humans, attributes such as ideology, dispositions, or values are better understood as reflections or downstream manifestations of latent preferences. However, language models do not possess preferences in the human sense. As next-token prediction machines, their behavior is shaped by statistical associations learned from human text corpora. This distinction allows us to invert the causal direction: rather than treating attributes as reflections of internal preference states, we tap into the pattern matching nature of LLMs and use attributes as control handles to steer model behavior.
To ensure that attributes effectively steer model preferences, we design a preset attribute bank that draws inspiration from theories across diverse social science disciplines, such as economics, political science, management science, and psychology. The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., values and beliefs, cognitive biases, or political ideology). These modes are further clustered into higher-level templates, which reflect different design purposes or theoretical orientations.
We define three primary templates. The core template contains general demographic descriptors and serves as the default baseline. The thematic template includes topic-specific modes such as economics, politics, culture and history. The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs Maslow (1943) and the Big Five personality traits John and Srivastava (1999). A condensed summary of these templates, representative modes, and attributes is shown in Table 1. These templates and modes are not fixed: they serve as a structured starting point but are fully extensible. Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts. By sampling and varying attributes from this flexible bank, we create prompts that instantiate distinct proxy agents occupying different regions of the latent preference space.
Template Representative Modes and Attributes
Core core: gender, race, ethnicity, education, income level,
Thematic economics: occupation, price sensitivity, wealth/debt,
politics: policy preferences, civic engagement, ideological orientation,
behavioral_biases: status quo bias, loss aversion, overconfidence bias,
Theoretical maslow: safety needs, self-actualization, love and belonging needs,
big_five: conscientiousness, neuroticism, extraversion,
Table 1: Structured view of the attribute bank, with representative modes and attributes grouped by template. Users can extend the bank with extra templates, modes, or attributes.
While the attribute bank provides structured coverage, we also enable dynamic attribute discovery. We introduce two additional freeform templates that are specific to the survey under analysis. The survey template derives attributes by prompting a designated LLM, called attribute learner, to identify relevant decision-making factors based on the overall survey topic covered by the training set of questions, $\mathcal{Q}_{\mathrm{train}}$. The question template goes further by asking the learner to derive attributes tailored to a specific survey question. In both cases, the inferred attributes are appended to the attribute bank only for the duration of the survey run. This hybrid strategy strikes a balance between theory-driven and data-driven endowment generations, and enables human-AI co-discovery of salient preference descriptors.
With the attribute bank in place, we build an endowment generator using a dedicated LLM termed endowment model that takes a set of attributes as input and produces diverse agent profiles by varying the instantiation of each selected attribute.
3.3 Active Endowment Generation
To construct a diverse preference basis without exhaustively enumerating attribute combinations, we create an active endowment generation procedure to operate under resource constraints. It proceeds in iterative steps. At each step, new endowments are sampled, instantiated as agents to elicit responses to survey questions, and evaluated based on their contribution to response diversity. The remainder of this section describes the key components of this procedure, including our variability scoring strategy, adaptive sampling steps, and entropy-based patching mechanisms.
Tracking response diversity with variability score
As established in Section 2, observed responses can serve as behavioral manifestations for latent preference vectors. Through this link, response variability serves as an initial diagnostic for assessing diversity in the underlying preference space. To effectively measure response variability, we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (1)
where $H_{i}(\mathcal{A})$ denotes the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ is the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ is the empirical proportion of responses selecting option $k$. The denominator normalizes for scale, ensuring that entropy is comparable across questions with different numbers of choices. A high question entropy indicates greater response diversity for that question. By varying the group $\mathcal{A}$, e.g., restricting it to agents sampled from a particular mode, we can compute question entropy at the group level and compare the diversity contributes across modes.
To turn question entropy into a useful metric for adaptive sampling, we introduce variability score for each mode. Formally, it is defined as
$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}.$ (2)
where $H_{i}(\mathcal{A}_{\mathrm{mode}})$ denotes the entropy of responses to question $i$ based on the cumulative set of agents from a given mode up to the current generation step, and $H_{i}(\mathcal{A}_{\mathrm{total}})$ is the corresponding question entropy across all agents generated so far. Computed this way, the variability score favors modes that are capable of generating endowments which elicit diverse responses to questions that other modes typically fail to diversify. The intuition is to maximize the informational utility of each observation under budget constraints. When most agents agree on a question, a divergent response signals a valuable distinction in the underlying preference space. In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round. The scores are then passed through a softmax function to produce a probabilty distribution over modes, which is used to guide endowment sampling in the next generation step.
Adaptive sampling and update steps
In P2P, endowment generation takes several steps. At the outset of the generation, an attribute learner is invoked to infer attributes from the survey based on the training questions, yielding the survey template on top of the existing templates in the attribute bank. The system then enters an initial sampling stage, drawing an equal number of endowments from each mode. The endowments are then used to instantiate agents to elicit responses to training questions, $\mathcal{Q}_{\mathrm{train}}$. The collective responses are then passed to Tracker for the computation of the variability scores and the sampling probabilities for the next update step. At the next update step, endowment budget allocation follows the multinomial distribution. A higher probability exploits the top-performing mode, while randomness allows exploration of its under-performing peers.
Patching low-entropy questions
In addition to the existing trade-off between exploitation and exploration, we adopt an additional boosting strategy, which we call question patching. At the end of each update step, as the Tracker computes the variability scores for each existing mode, it also keeps track of the questions with the lowest entropy values. When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$n$ entropy questions, where $n$ is user-defined. The split of the endowment budget between the active generation and boosting generation is governed by a user-specified ratio.
Figure 3: Entropy change across update steps for each question during active endowment generation for ATP W42. Each column represents a question. Green indicates an increase in entropy from the previous step, while red indicates a decline.
To avoid wasting resources on persistently low-entropy and to promote cross-pollination between modes, we additionally introduce a mixed mode strategy. It is triggered after a question appears $n$ times in the lowest-$k$ entropy list, where $n$ is a user-defined threshold. When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode. This encourages diversity injection from high-performing modes while maintaining a question-specific focus.
The adaptive sampling cycle continues until the generation budget $N_{A}$ (an upper bound on the maximum number of agents allowed) is exhausted. By iteratively targeting low-diversity questions and expanding the attribute space, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space.
3.4 Regression-Based Aggregation
With the candidate proxy basis assembled, P2P enters the second stage: Regression-based Aggregation. The goal at this stage is to find a functional basis from the candidate vectors to reconstruct the observed preference. Depending on the endowment budget and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables than observations. Consequently, a variable selection algorithm is vital. Besides, even with $N_{\mathrm{train}}+N_{\mathrm{valid}}>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations. To conduct variable selection, we consider regression methods with an added L1 penalization term. We can use either hold-out validation or cross-validation (CV) for hyperparameter tuning. The objective of the learning problem can be stated as
$\min_{\mathbf{w}}\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\sum_{k=1}^{K_{i}}\left(r_{ik}-\sum_{j=1}^{N_{A}}w_{j}d_{j,ik}\right)^{2}+\alpha\Omega(\mathbf{w}),$ (3)
$\mathrm{s.t.}\quad\sum_{k=1}^{K_{i}}\sum_{j=1}^{N_{A}}w_{j}d_{j,ik}=1\quad\forall i\in\mathcal{Q}_{\mathrm{train}},$ (4)
where $\mathbf{w}\in\mathbb{R}^{N_{A}}$ is the weight vector for the agents, $r_{ik}$ denotes the ground-truth proportion of responses selecting option $k$ for question $i$, $d_{j,ik}$ is a dummy response variable equal to $1$ if agent $j$ selects option $k$ and 0 otherwise, and $\alpha$ is the hyperparameter for the penalization. When $\Omega$ consists of only a L1 norm, the formulation amounts to constrained lasso. When $\Omega$ subsumes an additional L2 norm, it becomes constrained elastic net. Our simulation studies indicates that constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations (Appendix A.1), and it is potent in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix A.2). Both results corroborate using it as a selection backbone. Our simulation study also investigates the relationship between entropy and prediction fit—we leave the details to Appendix A.3 for interested readers.
4 Taking P2P to the Wild
As a first empirical test of the alignment framework, we deploy P2P on real-world social survey data. Specifically, we use the American Trends Panel (ATP), a nationally representative U.S. survey conducted by Pew Research Center in discrete waves. In this demonstration, we choose Wave 42, which probes trust in science by asking respondents about scientific competence, integrity, and public accountability. The wave includes responses from approximately 4464 U.S. adults Pew Research Center (2019) and the cleaned dataset contains 128 questions.
For the experiment, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation and test sets. For the endowment generation stage, we set the endowment budget to be $300$, with $10$ endowments generated for each mode during initial sampling. Subsequently, the rest of the endowments are generated across 5 update steps. In each step, we allocate 75% of the endowment budget to question patching, where the lowest 3 questions are considered. We set the trigger threshold of the mixed mode strategy at 1, so that the budget is reallocated if a question enters the lowest-3 list the second time. At the end of the generation, 5 questions are patched, with additionally 5 mixed modes created.
Endowment generation improves entropy
Figure 3 displays the entropy change for each question across the update steps. Evidently, a significant number of questions experience a noticeable rise in entropy at the end of the first update step. From the second step onward, mixed mode enters the generation loop. Marginal gains in entropy are observed for the majority of questions, whereas some experience mild entropy drops. Appendix B provides question entropy trajectories in levels. The three questions that exhibit the largest drops in entropy are among the highest-entropy questions at the initial sampling stage. The evolution elucidates the tradeoff made by the active endowment generator encoded in the variability scores.
For the selection of the agent ensemble, we fit both a constrained lasso and a constrained elastic net, using CV to select the penalization parameters. Consequently, lasso selects 53 out of the 300 synthetic agents to form the agent ensemble, while elastic net selects 43. The training and testing performances are similar between the two methods.
Figure 4: Prompts to Proxies: Emulation Results for ATP W42. Top: Train (left) and test (right) prediction performances for the binarized responses. Bottom: Human and agent (pre- and post- alignment) aggregated responses for the sample test question RQ4_F1Ae_W42.
Prediction indicates feasibility
Figure 4 shows the prediction performances of the agent ensemble selected with constrained lasso. The train-time root mean squared error (RMSE) is 0.08, with test RMSE slightly higher at 0.09, indicating adequate generalization ability of the functional basis. The bottom panel of Figure 4 presents a snapshot for the aggregated responses for a test set question, comparing the human ground truth with agentic emulations before and after regression-based aggregation. Overall, the results are encouraging, especially considering that P2P constructs the endowments based solely on preset and learned attributes and is agnostic on the ground-truth demographic data.
The panoramic view and future directions
In Appendix B, we provide a preliminary panel study, applying P2P to 14 ATP waves, each repeated 3 times. Performance variation is relatively low within each wave but considerably higher across the panel. In 8 out of 14 waves, P2P consistently achieves a test MSE below 0.015 across all repeats. Four waves exhibit test MSEs fluctuating around 0.015, while the remaining two fall in the range of 0.020 to 0.025. While our simulation studies suggest that entropy coverage, the number of observations and the size of the proxy pool all influence emulation performance, the preliminary panel results signal intriguing directions for future study. Understanding the sources of variation in empirical performance—particularly in relation to both the similarities and the distinctions from simulation results—offers a promising lens on the dynamics of human–AI alignment. These insights can also inform refinements to P2P ’s modular architecture and motivate extensions of its functionality.
P2P as a means to an important end
At a broader level, our framework operationalizes pluralistic alignment Sorensen et al. (2024). Its structured use of attribute-based control enables the representation of diverse value systems, while its regression-based weighting mechanism supports distributional pluralism without enforcing demographic fidelity. Crucially, P2P demonstrates that steerability and pluralism are not separate challenges but deeply intertwined: the ability to steer model behavior through promptable attributes is what enables meaningful diversity in aggregate preference reconstruction.
Although developed in the context of survey emulation, P2P generalizes naturally beyond surveys. The underlying preference reconstruction logic can be applied to other forms of observational data by discretizing continuous variables and binarizing categorical ones. Moreover, while this paper focuses on modeling aggregate population preferences, the framework could be adapted for individual-level alignment—for example, by learning proxy ensembles that emulate preferences of a single person across varied contexts.
Beyond its methodological contributions, P2P affords a modular platform for empirical alignment. Our work is among the first to streamline the use of large-scale social survey data—not as textual corpora for language model pretraining, but as a structured, theory-informed substrate for steering models toward specific populations, cultural milieus, and value contexts. It supports systematic evaluation of prompt engineering techniques, attribute design strategies and sampling heuristics. Researchers can test how different endowment structures affect response diversity, and how entropy metrics reflect or obscure latent preference variation. These tools offer a bridge between high-level alignment goals and grounded behavioral evaluation.
Finally, P2P offers a testbed for social scientists to probe and operationalize preference theories. By investigating how assumptions about cognitive style, value orientation, or belief structure affect response distributions and mode-level entropy, researchers can explore the explanatory power to theoretical constructs in a controlled but flexible setting. In this sense, the system fulfills the spirit of its name: P2P connects peer disciplines, translates prospects into products, carries ideas from pseudocode to paper, and moves pluralistic alignment toward practical needs.
5 Conclusion
This work establishes a theoretical foundation for viewing preference alignment as a problem of preference reconstruction. Inspired by revealed preference theory, our approach formalizes alignment as a two-stage process: first, constructing a functional basis of proxy agents using attribute-guided prompting; and then, recovering aggregation weights through supervised learning. In doing so, our work offers the first tractable implementation of pluralistic alignment, along with quantitative metrics to assess both diversity and fidelity.
Beyond its theoretical contributions, this framework offers a practical platform for interdisciplinary research. In the social sciences, it can support survey design, question testing, and nonresponse mitigation, serving as a cost-effective complement to traditional data collection. For researchers in LLM alignment, the modularity of our system provides a controlled setting to evaluate prompt engineering strategies and their influence on agent diversity and downstream performance. More broadly, it lays the groundwork for agentic systems that enable value-sensitive modeling across disciplines. Notably, P2P repositions social survey data not as training material, but as a guiding signal for culturally grounded model alignment.
Appendices
Appendix A Simulation Studies
In order to gauge the validity of the proposed preference model and the factors affecting alignment performance, we conduct three simulation studies where we have the privilege of ground-truth knowledge.
Simulation Design
For the simulation studies, we use the American Trends Panel (ATP) Wave 42 (W42) as the survey source, covering the topic of public trust in science and scientists. We run the active endowment generation on the survey to get 300 persona endowments, and partition the survey questions into training, validation and test sets by the 0.7:0.15:0.15 ratio. We subsequently simulate the responses to the survey questions using endowed GPT-4o agents as the synthetic respondents. The ATP W42 dataset has in total 129 questions, which after binarization leads to 517 observations—85% of these observations form the train-and-validation (trainval) set and the rest 15% is used to test the predictive accuracy of the selected agent ensemble.
Construction of the Ground Truth Data
To construct the ground truth data, we sample a fraction of the endowments as the ground truth and randomly assign weights to these endowments, which represent the proportions of the associated personae in the ground truth population. The ground truth data—percentages of individuals choosing difference options for a given question—are computed using weighted averages of responses provided by the ground-truth agents. The unselected endowments are used to construct the pool of proxy agents.
Regression Method
For the simulation studies, we choose constrained lasso as our main regression alignment method for investigation. As a secondary choice, constrained elastic net is an extension of constrained lasso that trades off between the L1 and L2 penalty terms. It is expected to have more stable performance when the features (agent responses) are highly correlated. Our initial simulations indicate that constrained elastic net demonstrates near identical performance to constrained Lasso, partly because the latter is a special case of the former and we are using cross-validation to select the best penalization parameters. Therefore, for the sake of building intuitions, we focus on constrained lasso in the simulations and expect the results to be largely indicative of constrained elastic net’s performance.
A.1 Recovering Ground Truth Agents with Constrained Lasso
As a first test of constrained lasso as a variable selection algorithm, we aim to evaluate its ability to recover ground-truth agents from a broader candidate pool. To render the selection task nontrivial, we set the prevalence of the ground-truth agents to 0.3—out of 10 selectable agents, only 3 contribute to the construction of the ground-truth signal. As the proxy agents dominate the agent pool, an adequate number of observations becomes essential for Lasso to reliably identify the true contributors. As a key metric, we define observation-to-agent ratio (OAR) to be the ratio of the number of training and validation observations to the number of selectable agents. By default, the OAR for our simulation baseline is 1.203.
To develop a panoramic view of constrained lasso’s ability in ground truth recovery, we conduct multiple simulation rounds under varying OARs. We generate different OAR variations using two complementary subsampling strategies:
  • Subsampling endowments. To increase the OAR, we adopt a subsampling strategy on endowments, where we subsample from the pool of selectable endowments, varying the total number of candidate agents while keeping the proportion of ground-truth agents fixed at 30%. The ground-truth weights are renormalized to sum to 1. We vary the subsample fraction from 0.5 to 1 in 20 equally spaced steps.
  • Subsampling questions. To decrease the OAR, we subsample survey questions while keeping the original train-validation-test split and ratio unchanged. This mimics a setting with reduced behavioral signal for inference. As in the endowment-based approach, we use 20 subsample fractions ranging from 0.5 to 1.
We repeat each subsampling procedure 10 times to construct the final dataset used in our simulation study. Figure 5 shows the trends of the mean squared error of the test set and the precision—that is, the proportion of selected agents who possess ground-truth endowments—across different OARs. To further illustrate the dynamics of selection, Figure 6 presents three representative snapshots of simulation runs, corresponding to low, medium, and high OAR settings.
As the figures indicate, constrained lasso’s ability to recover ground-truth agents from the candidate pool improves with a higher observation-to-agent ratio. This is intuitive as more observations offer a richer signal of both individual-level idiosyncrasies and aggregate preferences. With abundant behavioral cues, the selection algorithm is better equipped to distinguish true contributors from correlated siblings. In contrast, when observations are sparse, Lasso tends to relinquish its sharpness as a selection tool and instead defaults to identifying a functional basis that best explains the limited data. The mild decline in predictive accuracy under low OAR regimes suggests that this functional basis still approximates the aggregate preference reasonably well, albeit imperfectly.
Figure 5: Mean Squared Error (Test Set) and Precision by Different Observation-to-Agent Ratio (OAR) for Constrained Lasso. Each simulation is repeated 10 times with different random seeds. Subsampling of endowments is done with fixed ground-truth-to-proxy ratio. Subsampling of responses is done with fixed train-valid-test split and ratio.
Figure 6: Snapshots of Weight-vs-Coefficient Comparison and Prediction Performance. Each row represents a simulation run. The red bars on the left panel represent ground truth agent weights used to generate the aggregate data and the blue bars Lasso coefficients. The observation-to-agent ratios for the top, middle and bottom panels are: 0.617, 1.203, 2.407. Lasso’s ability to recover the ground truth agents from the agent pool degrades with lower observation-to-agent ratio, but its predictive accuracy only drops mildly.
A.2 Emulating Ground Truth Responses with Proxy Agents
While the first simulation exercise has readily allured to constrained lasso’s ability to find a functional basis to explain the observed data, in this second simulation study we extend the investigation by removing the ground-truth agents from the selectable agent pool. Under this setting, constrained lasso needs to fully rely on proxy agents to recover the ground-truth patterns. We analyze two distinct scenarios:
  • Subsampling proxies. We fix the total number of observations and reduce the proxy pool by retaining only a fraction as selectable agents. We sweep across the fractions from 0.05 to 1 in 20 equal spaced steps.
  • Subsampling training and validation questions. We fix the proxy pool but retain a fraction of training and validation observations. Similar to subsampling proxies, we construct 20 simulation rounds gradually raising the retained fraction from 0.05 to 1.
Like in Simulation Study 1, we repeat each simulation 10 times using different random seeds to form the final results. In Figure 7, the left panel displays the mean squared errors (MSE) and coefficients of determination ($R^{2}$) for the subsampling proxies scenario, while the right panel showcases the associated metrics for the subsampling training and validation questions scenario.
Figure 7: Left: Mean Squared Errors and Coefficients of Determination ($R^{2}$) by Number of Proxy Agents. Right: Mean Squared Errors and Coefficients of Determination ($R^{2}$) by Number of Observation. Each simulation is repeated 10 times with different random seeds.
Notably, constrained lasso exhibits robust performance across all simulated conditions. Even with only 10 selectable proxy agents, it achieves a mean test MSE of $0.0067$, yielding an $R^{2}$ of $0.93$. Predictive accuracy improves as we increase the number of selectable proxy agents, plateauing at $0.99$. In comparison, decreasing the number of observations has a lesser effect on predictive performance: in the recorded run with the lowest number of observations at 13, constrained lasso attains a test MSE of $0.0019$ and an $R^{2}$ of $0.98$. Figure 8 displays representative snapshots of the simulation runs.
Figure 8: Snapshots of Prediction Performance for Simulation Study 2.
The high predictive accuracy of the simulation runs indicates that constrained lasso can efficiently find a functional basis from a select few agents and determine the agent weights to emulate the ground-truth data patterns based on limited observations. This is in fact not surprising because in the initial assignment of the endowments into the ground truth and proxy groups we have used a random partition. As informed by our proposed behavioral preference model, this causes the ground truth vectors—used in the simulation study to form the aggregate preference—and proxy vectors to span the same preference space. As the preference space is not directly accessible, we can derive a partial gauge through the observed agent responses. If the responses of a group of agents have higher variability, it is an indication that the preference space they span is large. Conversely, if agents tend to agree on their responses to different questions, it is a signal that their associated vectors in the preference space might overlap in sub-dimensions, leading to a pool generalization ability.
To effectively measure response variability, in the paper we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (5)
where $H_{i}(\mathcal{A})$ is the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ denotes the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ represents the empirical proportion of responses selecting option $k$. The normalization factor ensures comparison across questions. A high question entropy indicates that the responses are highly varied for this question.
For the simulation studies, we define group entropy as the average question entropy computed using responses from that group:
${H}(\mathcal{A})=\frac{1}{N}\sum_{i=1}^{N}H_{i}(\mathcal{A}).$ (6)
The group entropy is calculated using responses to all survey questions, regardless of the split. Based on it, we introduce the notion of Entropy Coverage Ratio (ECR) defined as the ratio between the group entropy of proxy agents and that of the ground-truth agents:
$ECR=\frac{{H}(\mathcal{A}_{\text{proxy}})}{{H}(\mathcal{A}_{\text{gt}})}.$ (7)
An ECR lower than 1 indicates that the proxy agents have less response variability than the ground truth, while an ECR greater than 1 indicates the opposite.
Figure 9: Entropy Coverage Ratio Diagnostics for Simulation Study 2.
The left panel of Figure 9 shows the distribution of entropy coverage ratios (ECRs) for the two analyzed scenarios. The mass is concentrated around $1.0$, corroborating the earlier observation that random partitioning tends to preserve entropy structure. The right panel plots ECR against the number of proxies for the proxy-subsampling scenario. The relationship between ECR and the number of proxies is subtly positive: when the number of proxies is small, the likelihood of drawing an unrepresentative sample increases. As the sample size grows, group-level entropy converges toward the population-level entropy—an empirical manifestation of the law of large numbers.
A.3 Response Variability Matters
In this final simulation study, we analyze the effect of the entropy coverage ratio on predictive accuracy for constrained lasso.
Figure 10 displays the group entropies for THE W42 endowments organized by modes—we call them mode entropies in the endowment generation logic. We subsume them further into three different tiers according to the group entropy value: in total, we have 71 endowments in the low entropy tier, 92 in the mid entropy tier, and 137 in the high entropy tier.
Figure 10: Endowment Count per Mode (Sorted by Entropy).
To cover a wide range of ECR ratios, we design the following simulation strategy:
  1. 1.
    Ground Truth Construction. For each simulation round, pick one of the three tiers as the ground-truth tier and sample 30 endowments from the tier to construct the ground-truth agents and assign ground-truth weights. The unselected endowments from the tier are joined with the endowments from the other two tiers to form the pool of proxy candidates.
  2. 2.
    Proxy Agents Construction. For each simulation run, we select 30 proxy agents to form the proxy pool. We begin by selecting agents from the lowest entropy modes (ensuring that the cumulative candidate endowments surpass 30), and in each consecutive run, include the next mode with higher entropy into the selectable modes. We sweep across the modes by increasing mode entropy.
  3. 3.
    Proxy Agents Selection and Response Emulation. In each simulation run, after the proxy pool has been constructed, we use constrained lasso to select proxy agents to emulate the observed aggregate response data.
Figure 11: Mean Squared Error and Coefficient of Determination ($R^{2}$) by Entropy Coverage Ratio (ECR) for the trainval set (left) and test test (right). Entropy Coverage Ratio is defined as the ratio of the average response entropies of the proxies and ground-truth agents. LOWESS curves are fit using 10 repetitions.
We repeat each simulation 10 times to form the final results. Figure 11 visualizes the train and test performance of the constrained lasso by the entropy coverage ratio. As the ECR increases, the predictive accuracy improves at both the training and the test times, lending further credence to the analysis based on the preference model in the second simulation study.
Figure 12: Snapshots of Prediction Performance for Simulation Study 3.
Figure 12 offers three snapshots of the simulation runs corresponding to ECRs equaling $0.50$, $1.01$ and $1.67$. When proxy entropy is low compared with ground truth, the proxy agents fail to provide a sufficiently expressive basis to span the aggregate ground-truth preference. As a result, constrained Lasso can only approximate a lower-dimensional projection of the true preference signal. This results in poor prediction performance, as illustrated in the leftmost snapshot. Conversely, when proxy entropy exceeds ground-truth entropy, prediction error begins to increase—albeit at an infinitesimal scale. The subtle deterioration of performance when ECR surpasses 1 could be due to various reasons. A plausible cause is from the data perspective: when the proxy pool becomes too diverse, it may introduce spurious variability that does not align well with the true underlying structure. Put differently, an observed response might amalgamate signals from different latent preference factors, some of which are irrelevant to the aggregate preference in focus. While greater diversity expands the representational capacity of the proxy pool, it may also dilute the signal and induce overfitting, thereby reducing generalization accuracy.
The key takeaway from the simulation study is that absent the knowledge of ground truth, it is crucial to ensure an adequate group entropy among the proxy agents to form a functional basis for ground-truth emulation. However, one should also refrain from solely relying on entropy as a measure of preference diversity lest spurious correlations beguile us under limited data regimes. While our alignment method is intentionally designed to function with minimal data—requiring only aggregate response data—in practice, if individual human responses are available, a practitioner may use the group entropy computed from these responses as an anchor for endowment generation.
Appendix B Taking P2P to the Wild: Supplementary Experiment Results
B.1 Wave 42
Tracking entropy during active endowment generation
Figure 13 shows the entropy trajectories by question during active endowment generation. Of the tracked questions, 53 experience a clear rising trend in entropy, while 3 see a noticeable fall in entropy at the end of the generation, the other questions have stable entropy values across the update steps. Figure 14 displays the question entropies at the end of active endowment generation.
Figure 13: Entropy trajectories by question during active endowment generation. Trends are classified based on slope and volatility of the entropy trajectory: “rising” if the slope exceeds 0.01 and standard deviation is above 0.02; “falling” if the slope is below –0.01 with sufficient volatility; otherwise labeled “stable”.
Figure 14: Question entropies at the end of active endowment generation.
Lasso Selection
Figure 15 plots the cross-validated mean squared error (MSE) for each value of $\alpha$ examined during hyperparameter tuning. Among the 30 candidate values, cross-validation identifies the optimal $\alpha$ as $1.17\times 10^{-3}$. Figure 16 shows the number of agents selected by constrained lasso by mode entropy, together with the number of agents selected from each template/strategy. Mixed modes contribute the largest share to the ensemble with 20 proxy agents, followed by question-patching modes with 14. Thematic modes come next, contributing 5 agents, alongside 5 from the survey mode. The remaining 1 proxy is selected from the core mode.
Figure 15: Cross-validated MSE by alpha ($\alpha$), the penalization parameter, for constrained lasso. The best alpha selected by CV is $1.17\times 10^{-3}$.
Figure 16: Number of agents selected by constrained lasso by mode entropy. Mixed modes contribute 20 proxies to the ensemble, followed by question patching modes contributing 14. The thematic modes come third, contributing 5 proxies. Five proxy agents are selected from the survey mode, while the rest 1 is from the core mode.
Elastic Net Results
In the main paper, we show prediction results using constrained lasso. Here, we provide the results for constrained elastic net (Figures 17 and 18). Overall, the prediction results are similar those using constrained lasso.
Figure 17: Cross-validated MSE by alpha, the overall penalization parameter, and L1 ratio, the penalization weight for the L1 term, for constrained elastic net. The lowest cross-validated MSE is achieved at $\alpha=2.59\times 10^{-3}$, with l1 ratio set at $0.60$.
Figure 18: Prediction performance of the agent ensemble selected by constrained elastic net. Results are similar to those using constrained lasso.
B.2 Panel Study Across ATP Waves
To obtain a panoramic view of P2P’s performance empirically, we conduct a preliminary panel study applying P2P to a collection of 14 waves from the American Trends Panel (ATP). Each wave is repeated three times to assess the consistency and robustness of our method under varying random seeds and endowment draws. This subsection provides an extended empirical perspective on the real-world applicability of our approach and sheds light on variability patterns that may not arise in controlled simulation environments. The ATP waves used in this panel study are: W26, W27, W29, W32, W34, W36, W41, W42, W45, W49, W50, W54, W82, and W92.
Figure 19 illustrates the test MSEs for the 42 experiments organized by the number of survey questions per wave. As shown, within-wave performance is generally stable, with most waves exhibiting low variance in test MSE across repeats. However, test MSE varies more significantly across waves. The performance distribution is summarized below:
  • 8/14 waves achieve test MSE consistently below 0.015.
  • 4/14 waves fluctuate around 0.015.
  • 2/14 waves have higher error ranges (0.020–0.025).
These preliminary results reveal no clear relationship between test MSE and the number of survey questions, suggesting that wave-specific factors, such as survey content, attribute coverage, or question formulation, may play a larger role in shaping performance.
Figure 19: Test performance across ATP waves. Each point represents a wave-specific run, colored by wave identity. The horizontal dashed lines mark two threshold levels: 0.015 and 0.020, used to classify performance regimes.
To illustrate the interaction between question-level entropy dynamics and predictive performance, we include three snapshots of the panel study in Figure 20, showcasing the differences in question entropy trajectories during active endowment generation and prediction performance.
Figure 20: Representative snapshots for three ATP waves with different test-time MSE. Top: the worst-performing wave. Middle: a moderately performing wave. Bottom: the best-performing wave. Left: entropy trajectories classified as rising (green), falling (red), or stable (gray). Center/right: Lasso regression fits for aggregate response prediction. While rising entropy often coincides with improved fit, some waves achieve strong performance with limited entropy change, suggesting additional contributing factors beyond diversity expansion.
Interestingly, the wave with the lowest test MSE (bottom row) exhibits relatively modest entropy change, while the poorest-performing wave (top row) shows mostly flat entropy evolutions. This suggests that while entropy dynamics can be indicative of diversity expansion, strong initial endowment coverage or other wave-specific factors (e.g., question content or noise) may also influence alignment outcomes. Entropy trends are a useful diagnostic but should be interpreted in the context of wave-specific factors.
The patterns observed in the panel study warrant deeper analysis, which is likely to inform refinements that make P2P more stable and effective across settings. One especially promising direction is to extend the entropy-based active endowment generation strategy to support multi-objective criteria, allowing the system to balance diversity, coverage, and alignment more systematically.
Another promising extension is to include a module that directly measures question semantics. Based on our preference reconstruction theory, survey questions define the response space $\mathcal{R}$. Question entropy serves as a key signal for underlying diversity when conditioning on a fixed question. However, an equally important source of variation lies in the diversity across questions themselves. This includes differences in topic coverage, difficulty, and framing—factors that shape the expressiveness and completeness of the response space. Ideally, the question set should span a semantically broad topic space so that variations in response entropy meaningfully reflect differences in latent preferences. A dedicated semantic module would enable content-aware data-splitting strategies and support more rigorous generalization analysis across question types. This is an important direction for future development of P2P.
Due to budget constraints, our empirical evaluation relies on American Trends Panel (ATP) waves, which are topical in nature and tend to focus on a narrow set of issues per wave. While this allows for controlled testing, it limits the diversity of question types and themes. A more expansive and promising application domain for P2P is large-scale general social surveys, such as the General Social Survey (GSS) or World Values Survey, which span a wide range of topics and include high-quality, theoretically grounded questions relevant across disciplines. Testing P2P in such settings would provide a more comprehensive benchmark for its alignment and coverage capabilities, and yield better insights into its generalizability across domains of social inquiry. Moreover, the agent ensembles generated from general social surveys could, in turn, be used to provide feedback on smaller topical surveys to improve survey design and assess content quality.
Appendix C Methodological and Experimental Details
C.1 Prompts
C.1.1 Attribute Learner (AttributeLearner)
To generate attributes from a single question:
You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond.
You are given a single survey question. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question.
Focus on underlying factors that would cause meaningful variation in answers across different types of people. Avoid generic or overly broad attributes.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["religious affiliation", "political ideology", "trust in government"]
To generate attributes from a set of questions:
You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond.
You are given a set of training-only survey questions. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions.
Carefully analyze the content and framing of the questions. Identify underlying factors that might shape how different people respond. Focus on attributes that are salient, discriminative, and potentially variable across respondents.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
C.1.2 Endowment Model (EndowmentModel)
To generate endowments from attributes:
You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys.
Each persona should include:
‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female). No punctuation or spaces.
‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent.
Instructions:
Represent a wide range of age, gender, race, education, region, and political ideology.
Avoid repetition of phrasing or demographic combinations across personas.
Do not include explanations or formatting outside of the JSON array.
C.1.3 Survey Conductor (SurveyConductor)
To use an agent to answer a survey question:
You are completing a survey.
Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences.
Respond with only the final answer string, not the code or label in brackets.
Do not include any reasoning, explanation, or commentary.
Do not preface your answer with phrases like ’I would choose’.
Just return the answer text exactly as it appears in the options.
C.2 Active Endowment Generation
Algorithm 1 Active Endowment Generation
1: Initialize attribute bank $\mathcal{B}$ with core, thematic, and theoretical templates
2: Derive survey-specific attributes from $\mathcal{Q}_{\mathrm{train}}$ via AttributeLearner; append to $\mathcal{B}$ Initial Sampling Stage
3: Sample initial endowments $\mathcal{E}_{0}$ via equal-mode sampling from $\mathcal{B}$
4: Instantiate initial agents $\mathcal{A}_{0}$ from $\mathcal{E}_{0}$ and elicit responses $\mathcal{R}_{0}$ on $\mathcal{Q}_{\mathrm{train}}$
5: Initialize ThemeVariabilityTracker (henceforce, Tracker) with $\mathcal{A}_{0}$, $\mathcal{R}_{0}$, and $\mathcal{Q}_{\mathrm{train}}$. Expansion Stage
6:while agent budget $N_{A}$ not yet reached do
7:  Compute variability scores and mode sampling probabilities via Tracker
8:  Allocate endowment budget across modes according to probabilities
9:  if question patching is enabled then
10:   Identify lowest-entropy questions $\mathcal{Q}_{\mathrm{low}}$ via Tracker
11:   for each $q\in\mathcal{Q}_{\mathrm{low}}$ do
12:    if $q$ is queried for the first time then
13:     Extract attributes for $q$ via AttributeLearner and append to $\mathcal{B}$
14:    else if $q$ has appeared more than $n$ times then
15:     Enable mixed-mode strategy using $q$ and top-performing mode
16:    end if
17:   end for
18:  end if
19:  Generate new endowments $\mathcal{E}_{\mathrm{new}}$ using allocated budget over sampled modes, patched question modes (if any) and mixed modes (if any).
20:  Instantiate agents $\mathcal{A}_{\mathrm{new}}$ and elicit responses $\mathcal{R}_{\mathrm{new}}$ on $\mathcal{Q}_{\mathrm{train}}$
21:  Update: $\mathcal{A}\leftarrow\mathcal{A}\cup\mathcal{A}_{\mathrm{new}}$, $\mathcal{R}_{\mathcal{A}}\leftarrow\mathcal{R}_{\mathcal{A}}\cup\mathcal{R}_{\mathrm{new}}$
22:  Update Tracker with $\mathcal{A}$, $\mathcal{R}_{\mathcal{A}}$, and $\mathcal{Q}_{\mathrm{train}}$
23:end while
24:Return: Final agent pool $\mathcal{A}$ and full response set $\mathcal{R}_{\mathcal{A}}$
Algorithm 2 Tracker Update Procedure
1: Input: Current agent pool $\mathcal{A}$, current agent responses $\mathcal{R}_{\mathcal{A}}$, and training questions $\mathcal{Q}_{\mathrm{train}}$
2:for each question $i\in\mathcal{Q}_{\mathrm{train}}$ do
3:  Compute question entropy $H_{i}(\mathcal{A})$
4:end for
5:for each mode do
6:  Compute variability score $V(\mathcal{A}_{\mathrm{mode}})$
7:end for
8: Compute softmax probabilities from variability scores
9: Output: Variability scores, question entropies, and sampling probabilities.
C.3 Example Endowments
An example of an endowment for a progressive female in her 20s living in an urban area
[…]
A 24-year-old urban female who identifies as bisexual and holds a master’s degree in social work. She exhibits a high fairness bias and is often concerned about equality. Despite her awareness, she sometimes struggles with extreme response bias, influenced by her strong convictions.
An example of an endowment for a nonbinary in their 30s valuing liberty living in a suburban area
[…]
They are a 32-year-old nonbinary software developer living in the suburbs. They highly value personal liberty and have a strong work ethic, often putting in extra hours for their projects. With a master’s degree in computer science, they frequently engage with scientific research, feeling that understanding complex topics fosters innovation. Their income is above average, and they embrace intelligence as a vital asset for personal growth.
An example of an endowment for a white male progressive in his 60s living in a small town
[…]
Pat is a 67-year-old White male from a small town. He identifies as a progressive and believes strongly in the importance of intelligence in community discussions and personal growth. Pat, who identifies as heterosexual, is skeptical of the increasing influence of big corporations on local politics, and he often seeks media that highlights grassroots movements and community activism. He balances his media consumption between traditional outlets and local storytelling platforms, knowing the value of good journalism.
Appendix D Extended Literature Review
D.1 Large Language Models to Emulate Human Preferences
Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization Tseng et al. [2024] has been largely focused on emulating human responses using various techniques. These include learning from feedback or reward models Bai et al. [2022], Christiano et al. [2017], under distribution shifts Leike et al. [2018], Krueger et al. [2021], finetuning Tan et al. [2024], retrieval augmented generation Prahlad et al. [2025], representation learning Ren et al. [2024], and using human-generated Horton [2023], Castricato et al. [2025], Argyle et al. [2023] prompts. However, when diverse responses need to be generated from many users or user groups instead of individuals, the data and time resources required for alignment grow linearly with the number of LLMs that need to be aligned. To reduce alignment costs, some works instead rely on AI-generated Schuller et al. [2024], Salminen et al. [2024], Ge et al. [2025], Simmons [2023] prompts, or alignment to multiple user groups Zhao et al. [2024], Mondal et al. [2025], even entire populations Cao et al. [2025]. However, even the best models struggle with this task Cao et al. [2025]. Furthermore, responses from LLMs have been found to have less variation than responses from real surveys Bisbee et al. [2024], and LLMs have a tendency to respond in the middle category Wang et al. [2024]. Consequently, it is imperative to address these limitations and devise data- and time-efficient approaches that enable LLMs to more accurately and reliably serve as proxies for survey populations. Our work contributes to this goal by introducing an active learning–inspired alignment framework that dynamically constructs a diverse agent basis under resource constraints and reconstructs population-level preferences via regression-based aggregation, without relying on ground-truth demographic profiles.
D.2 Data Selection and Active Learning
To increase preference coverage while operating under budget constraints, our active endowment generation pipeline draws inspiration from machine learning paradigms, particularly active learning.
Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization Coleman et al. [2020], Paul et al. [2021]. Selection via Proxy (SVP) Coleman et al. [2020] uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores Paul et al. [2021] identify important training examples early in training.
Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning Astorga et al. [2024], Hübotter et al. [2025], Li et al. [2025]. Active learning strategies balance exploitation, focusing on uncertain or high-impact samples, with exploration, seeking diverse and novel samples to broaden coverage.
In our system, we apply this logic to guide both endowment generation and question patching. Specifically, low-entropy responses are treated as indicators of low coverage or model uncertainty. We refer to question patching as a boosting strategy—not in the ensemble learning sense, but to distinguish it from the main adaptive sampling process—since it targets underperforming questions by allocating additional sampling to enhance expressivity. Meanwhile, the exploitation–exploration mechanism used to sample across existing modes reflects the same active learning principle, but with a different emphasis: prioritizing high-utility regions while maintaining representational diversity. Together, these components form an active learning–inspired loop that incrementally expands coverage of the preference space while minimizing redundant agent generation.
Appendix E Mathematical Formalization of the Preference Reconstruction Theory
E.1 Space Definitions and Mapping Framework
We define four key spaces for formalizaing our alignment framework:
  • Attribute space $\mathcal{B}\subset\mathbb{R}^{d}$: each $b\in\mathcal{B}$ is a structured configuration of interpretable attributes drawn from the attribute bank (e.g., ideology, altruism, personality). Attributes serve as control handles that steer LLM behavior.
  • Endowment space $\mathcal{E}$: each $e\in\mathcal{E}$ is a natural language prompt or agent description instantiated from an attribute configuration $b\in\mathcal{B}$. Endowments serve as the actual inputs to the LLM. We treat $\mathcal{E}$ as a structured set, not a Euclidean space.
  • Preference space $\mathcal{P}\subset\mathbb{R}^{m}$: each $p\in\mathcal{P}$ represents a latent vector encoding a behavioral inclination or preference profile. This space is unobservable but assumed to mediate how attributes shape responses.
  • Response space $\mathcal{R}\subset\mathbb{R}^{n}$: each $r\in\mathcal{R}$ is an observable vector of responses to survey questions (e.g., binary or Likert-coded).
Based on the four spaces, we define three mappings:
  • Attribute-to-Endowment Mapping ($\phi:\mathcal{B}\to\mathcal{E}$): This is a stochastic mapping that transforms a symbolic attribute configuration into a textual prompt or persona instantiation readily parsable by an LLM. Since multiple distinct prompts can be generated from the same attributes, we model $\phi$ as sampling from a conditional distribution over prompts given attributes.
  • Endowment-to-Preference Mapping ($g:\mathcal{E}\to\mathcal{P}$): This captures the internal LLM process by which a given prompt (endowment) induces a latent preference state. This mapping is governed by the LLM’s internal representations, shaped jointly by its large-scale pretraining corpora, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF). Albeit unobservable, this preference vector mediates the model’s behavioral tendencies in downstream response generation.
  • Preference-to-Response Mapping ($h:\mathcal{P}\to\mathcal{R}$): This final mapping generates a survey response vector based on the latent preference state. It represents how an agent’s inclinations are expressed as answers to survey questions.
E.2 Preference Reconstruction via a Functional Basis
The overall agent behavior function is thus a composition of the three mappings:
$f=h\circ g\circ\phi:\mathcal{B}\rightarrow\mathcal{E}\rightarrow\mathcal{P}\rightarrow\mathcal{R}$
where $f(b)$ denotes the response behavior of an LLM agent endowed based on attributes $b$.
Given a set of sampled agents $\mathcal{A}=\{b_{1},\dots,b_{k}\}$, we obtain a response basis $\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}$. We then treat alignment as a projection task:
$\min_{w\in\mathbb{R}^{|\mathcal{A}|}}\left\|\sum_{b\in\mathcal{A}}w_{b}f(b)-r_{\text{pop}}\right\|^{2}$ (8)
where $r_{\text{pop}}$ is the target aggregate response vector, and the weights $w_{b}$ reflect the inferred contribution of each proxy agent to the population preference.
E.3 Alignment as a Two-Stage Problem
We frame the alignment process as a two-stage problem grounded in mappings across attribute, endowment, preference and responses spaces.
Stage 1: Constructing a functional basis based on response variability.
The goal of the first stage is to construct a diverse and expressive set of LLM agents whose latent preferences collectively span the space where the ground-truth population preference resides. Since latent preferences are not directly observable, we use response variability as an indirect gauge, leveraging the preference-to-response mapping ($h:\mathcal{P}\to\mathcal{R}$). Starting from a sampled set of attribute configurations $\mathcal{A}=\{b_{1},\dots,b_{k}\}\subset\mathcal{B}$, we apply the composition mapping $f:\mathcal{B}\to\mathcal{R}$ to obtain a set of simulated agent responses:
$\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}\subset\mathcal{R}.$
We interpret variability in $\hat{\mathcal{R}}$ as an indirect signal of the latent variation in $\hat{\mathcal{P}}$. The design objective is to ensure that the span of $\hat{\mathcal{R}}$ is sufficiently rich so that the latent preference vectors form a functional basis for reconstructing population-level behavior. To optimize this objective, we operate in the more interpretable and tractable attribute space $\mathcal{B}$ rather than the high-dimensional preference space $\mathcal{P}$. By actively exploring semantically meaningful regions of $\mathcal{B}$—analogous to domain-aware active learning strategies Hao et al. [2024] and latent-manifold transformations Jin et al. [2024]—we adaptively improve coverage of $\hat{\mathcal{P}}$.
Stage 2: Projecting onto population responses.
In the second stage, we assume access to a population-level response vector $r_{\text{pop}}\in\mathcal{R}$, such as aggregate survey results. The construction of the agent ensemble is done by solving the optimization problem defined by (8), where each weight $w_{b}$ reflects the inferred contribution of the agent defined by attribute set $b$ in reconstructing the aggregate signal.
Together, these two stages—constructing a coverage-maximizing basis and estimating population-aligned weights—enable efficient and demographically-agnostic preference reconstruction using steerable LLM agents.